OpenStudy (anonymous):

find the exact value of each of the remaining trigonometric functions of theta sin= -12/13,180degrees

OpenStudy (anonymous):

draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose sine (opposite over hypotenuse) is \(\frac{12}{13}\) what you need is the third side, which you find via pythagoras

OpenStudy (anonymous):

or else just remembering the famous \(5-12-13\)right triangle if you compute it , you would compute \[\sqrt{13^2-12^2}=\sqrt{136-144}=\sqrt{25}=5\]

OpenStudy (anonymous):

now you can take any trig ratio you want

OpenStudy (anonymous):



OpenStudy (anonymous):

just remember you are in quadrant 3 so both sine and cosine are negative

OpenStudy (anonymous):

so cos a is 5/13 is -.385?

OpenStudy (anonymous):

i meant cos theta

OpenStudy (anonymous):

minus since you are in quad 3 leave the decimal out of it, just \[\cos(\theta)=-\frac{5}{13}\]

OpenStudy (anonymous):

sso tan is 12/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and cot is 5/12 sec 13/5 csc 13/12?? do any of those need to be negative

OpenStudy (anonymous):

secant and cosecant are negative

OpenStudy (anonymous):

they are the reciprocals of sine and cosine both of those are negative in quadrant 3

OpenStudy (anonymous):

okay thanks !!!:))

OpenStudy (anonymous):

is cos-45 degrees -sqrt2/2